OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Use the following System of Equations to answer the following question: 2x + 5y = 15 –4x + 7y = –13 What can you say about the graphs of each equation? a. Both equations have a positive slope. b. The first equation has a negative slope while the second equation has a positive slope. c. The first equation is a parabola while the second is linear. d. The equations will form perpendicular lines. *not c*

OpenStudy (igreen):

Obviously it's not C..both are linear.

OpenStudy (igreen):

The rest compare the slopes..so we should change both of them into slope-intercept form. 2x + 5y = 15 Subtract 2x to both sides: 5y = -2x + 15 Divide 5 to both sides: y = -2/5x + 3 ------------------------ -4x + 7y = -13 Add 4x to both sides: 7y = 4x - 13 Divide 7 to both sides: y = 4/7x - 13/7 So the slope of the first one is -2/5, and the slope of the 2nd is 4/7.

OpenStudy (igreen):

So what do you think the answer is?

OpenStudy (anonymous):

@iGreen B ?

OpenStudy (igreen):

You got it.